A turboprop plane flying with the wind flew 1,400 mi in 5 h. Flying against the wind, the plane required 7 h to travel the same distance. Find the rate of the wind and the rate of the plane in calm air. Answer by galactus(183) 
 With the wind, we add the two rates. Against the wind, we subtract them.
Let rp=rate of plane and rw = the rate of thr wind.
With the wind:
Since d=rt, we have 1400=5(rp+rw)
Against the wind:
we have 1400=7(rp-rw)
These reduce to
rp+rw=280
rp-rw=200
Now, from the first equation we have rp=280-rw
Sub into the second equation:
280-rw-rw=200
280-2rw=200
rw=40=rate of wind.
That means the rate of the plane, rp, is 240.